OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)?

OpenStudy (anonymous):

can you give me the slope that passes through these two points? use m = (y1-y2)/(x1-x2) to find slope.

OpenStudy (anonymous):

(y-y1)/(y1-y2)=(x-x1)(x1-x2)

OpenStudy (anonymous):

@Dpalnc 0?

OpenStudy (anonymous):

once you find slope m, put it in point slope form: y - y1 = m(x-x1) then to get it in general form, make 1 side of the equal sign equal to 0.

OpenStudy (anonymous):

what is y?

OpenStudy (anonymous):

ok.. agreed. now that the slope is 0. notice that no matter what x is your slope is zero..

OpenStudy (anonymous):

so your equation is y = 2

OpenStudy (anonymous):

to make it in general form, subtract 2 from both sides to get y - 2 = 0